# Statistical Inference for Temporal Difference Learning with Linear Function Approximation

### v1

@@ A.1 Useful algebriac facts
$I_{2}$ $=\underset{I_{21}}{\underbrace{\frac{1}{T}\sum_{t=1}^{T}\bm{S}_{t}\bm{\Gamma}\bm{S}_{t}^{\top}}}+\underset{I_{22}}{\underbrace{\frac{1}{T}\sum_{t=1}^{T}\prod_{k=t}^{T}(\bm{I}-\eta_{k}\bm{A})\bm{\Gamma}\prod_{k=t}^{T}(\bm{I}-\eta_{k}\bm{A}^{\top})}}$
($-\underset{I_{23}}{\underbrace{\left(\frac{1}{T}\sum_{t=1}^{T}\bm{S}_{t}\bm{\Gamma}\prod_{k=t}^{T}(\bm{I}-\eta_{k}\bm{A}^{\top})+\frac{1}{T}\sum_{t=1}^{T}\prod_{k=t}^{T}(\bm{I}-\eta_{k}\bm{A})\bm{\Gamma}\bm{S}_{t}^{\top}\right)}}.$
Since $\|\bm{S}_{t}\|=O(T^{\alpha-1})$, it is easy to verify that $I_{21}=\bm{O}(T^{2\alpha-2})$; meanwhile, theorem A.7 indicates that $I_{22}=T^{\alpha-1}\bm{X}(\bm{\Lambda}^{\star})+O(T^{2\alpha-2})$. As for the term $I_{23}$, we observe that
$\left\|\sum_{t=1}^{T}\bm{S}_{t}\bm{\Gamma}\prod_{k=t}^{T}(\bm{I}-\eta_{k}\bm{A}^{\top})\right\|$ $\lesssim\sum_{t=1}^{T}t^{\alpha-1}\prod_{k=t}^{T}(1-\beta k^{-\alpha})$
$<T^{2\alpha-1}\sum_{t=1}^{T}t^{\alpha-1}\prod_{k=t+1}^{T}(1-\beta k^{-\alpha})$
$=T^{2\alpha-1}\cdot\frac{1}{\beta}\sum_{t=1}^{T}\left(\prod_{k=t+1}^{T}-\prod_{k=t}^{T}\right)(1-\beta k^{-\alpha})$
$<\frac{T^{2\alpha-1}}{\beta}.$
Consequently, it can be guaranteed that $I_{23}=\bm{O}(T^{2\alpha-2})$. In combintation, we have shown that
$T^{\alpha-1}\bm{X}(\bm{\Lambda}^{\star})-\bm{O}(T^{2\alpha-2})\preceq\frac{1}{T}\sum_{t=1}^{T}(\bm{Q}_{t}-\bm{A}^{-1})\bm{\Gamma}(\bm{Q}_{t}-\bm{A}^{-1})^{\top}\preceq T^{\alpha-1}\bm{X}(\bm{\Lambda}^{\star})+\bm{O}(T^{2\alpha-2});$
the theorem follows naturally by combining this with the bound on $I_{1}$.∎
@@
With $\bm{A}$ defined as (6a) and $\overline{\bm{\Lambda}}_{T}$ defined as in (18), when
@@ A.2.1 Berry-Esseen bounds
The following theorem, proved by Shao and Zhang, (2021), gives a Berry-Esseen bound for non-linear statistics which would be crucial for our analysis of the non-asymptotic distribution of the TD estimation error.
@@
(Theorem 2.1 of Shao and Zhang, (2021)) Let $\bm{X}_{1},...,\bm{X}_{n}$ be independent random variables, and let $\bm{T}(\bm{X}_{1},...,\bm{X}_{n})$ be a $d$-dimensional statistic that can be represented as
@@ A.2.2 Properties of multivariate Gaussian distributions
The following theorem proposed by Devroye et al., (2018) gives an upper bound for the total variation (TV) distance between two Gaussian random variables with same means and different covariance matrixes.
Theorem A.12 (Devroye et al., (2018)).
Let $\bm{\Lambda}_{1},\bm{\Lambda}_{2}\in\mathbb{S}^{d\times d}$ be two positive definite matrices, and $\bm{\mu}$ be any vector in $\mathbb{R}^{d}$. Then the TV distance between Gaussian distirbutions $\mathcal{N}(\bm{\mu},\bm{\Lambda}_{1})$ and $\mathcal{N}(\bm{\mu},\bm{\Lambda}_{2})$ is bounded by
@@
In the case of $\bm{\Lambda}_{1}\succ\bm{\Lambda}_{2}$, the following proposition further specifies that the TV distance and the distance on convex sets are the same bewteen $\mathcal{N}(\bm{0},\bm{\Lambda}_{1})$ and $\mathcal{N}(\bm{0},\bm{\Lambda}_{2})$.
@@
$\frac{p_{1}(\bm{x})}{p_{2}(\bm{x})}$ $=\frac{\frac{1}{\sqrt{\mathsf{det}(\bm{\Lambda_{1}})}}\exp(-\bm{x}^{\top}\bm{\Lambda_{1}}^{-1}\bm{x})}{\frac{1}{\sqrt{\mathsf{det}(\bm{\Lambda_{1}})}}\exp(-\bm{x}^{\top}\bm{\Lambda_{2}}^{-1}\bm{x})}$
@@
Combining these two inequalitites completes the proof. ∎
@@ A.3 Concentration inequalities
The following Azuma’s inequality for matrix martingales is shown in Tropp, (2011).
@@
Specifically, for vector-valued martingales, the following proposition is shown by Kallenberg and Sztencel, (1991).
@@ Appendix B Proof of main results
B.1 High-probability convergence guarantee for the original TD estimation error
In order to derive the high-probability convergence guarantee for $\overline{\bm{\Delta}}_{T}$, we firstly propose the following result on the time-uniform convergence of the original TD estimator $\bm{\Delta}_{t}$, under condition 12.
@@ B.1 High-probability convergence guarantee for the original TD estimation error
Consider the TD algorithm (9) with independent samples and decaying stepsizes as in condition 12. For any $\delta\in(0,1)$, there exists $\eta_{0}>0$, such that with probability at least $1-\delta$, the norm of $\bm{\Delta}_{t}$ is bounded by
@@
For simplicity, in this proof we deonte
@@
For any given $\delta$, we define $t^{\star}$ as a function of $\delta$, speicifically
@@
$\mathcal{H}_{i}=\{\max_{1\leq j\leq i}\|\bm{\Delta}_{j}\|\leq R\},\quad\text{and}\quad\widetilde{\bm{\zeta}}_{i}=\bm{\zeta}_{i}\mathds{1}(\mathcal{H}_{i-1}).$
@@
$\|\widetilde{\bm{\zeta}}_{i}\|_{2}\leq 4R+2\|\bm{\theta}^{\star}\|_{2}+1.$
@@
$\left\|\sum_{i=1}^{t}\bm{R}_{i}^{t}\widetilde{\bm{\zeta}}_{i}\right\|_{2}^{2}$ $\leq\sum_{i=1}^{t}\|\bm{R}_{i}^{t}\|^{2}\|\widetilde{\bm{\zeta}}_{i}\|_{2}^{2}$
@@
$\left\|\sum_{i=1}^{t}\bm{R}_{i}^{t}\widetilde{\bm{\zeta}}_{i}\right\|_{2}$ $<2\eta_{0}(4R+2\|\bm{\theta}^{\star}\|_{2}+1)\sqrt{\frac{2}{2\alpha-1}\log\frac{9t^{\star}}{\delta}}$
@@
$=\left\|\prod_{k=1}^{t}(\bm{I}-\eta_{k}\bm{A})\bm{\Delta}_{0}\right\|_{2}+\left\|\sum_{i=1}^{t}\bm{R}_{i}^{t}\widetilde{\bm{\zeta}}_{i}\right\|_{2}$
@@
$\left\|\sum_{i=1}^{t}\bm{R}_{i}^{t}\widetilde{\bm{\zeta}}_{i}\right\|_{2}$ $\leq 2\eta_{0}(4R+2\|\bm{\theta}^{\star}\|_{2}+1)\sqrt{2W_{t}\log\frac{3}{\delta_{t}}}$
@@
$\left\|\sum_{i=1}^{t}\bm{R}_{i}^{t}\widetilde{\bm{\zeta}}_{i}\right\|_{2}\leq 2\eta_{0}\left(\frac{2\alpha}{\beta}\right)^{\frac{\alpha}{2(1-\alpha)}}\sqrt{\frac{2}{2\alpha-1}\log\frac{9T}{\delta}}(4R+2\|\bm{\theta}^{\star}\|_{2}+1)t^{-\frac{\alpha}{2}}.$
@@
B.2 Proof of Theorem 3.1
Similar to the proof of Theorem B.1, we firstly expand $\overline{\bm{\Delta}}_{T}$ into three terms: a non-random term capturing the error introduced by the intialization $\bm{\theta}_{0}$, a martingale, and a sum of $i.i.d.$ vectors. Specifically, (B.1) implies
@@
The term $I_{2}$ is a sum of $i.i.d.$ random vectors, and can hence be bounded by invoking the matrix Berstein’s inequality. Specifically, define
@@
$\left\|\frac{1}{T}\sum_{i=1}^{T}\bm{Q}_{i}(\bm{A}_{i}\bm{\theta}^{\star}-\bm{b}_{i})\right\|_{2}\leq 2\sqrt{W_{\max}\log\frac{2d}{\delta}}+\frac{4}{3}B_{\max}\log\frac{2d}{\delta}.$
@@
$\left\|\frac{1}{T}\sum_{i=1}^{T}\bm{Q}_{i}(\bm{A}_{i}\bm{\theta}^{\star}-\bm{b}_{i})\right\|_{2}\leq 2\sqrt{\frac{\mathsf{Tr}(\overline{\bm{\Lambda}}_{T})}{T}\log\frac{6d}{\delta}}+\frac{4\eta_{0}}{T}\left(\frac{2}{\beta}\right)^{\frac{1}{1-\alpha}}(2\|\bm{\theta}^{\star}\|_{2}+1)\log\frac{6d}{\delta}.$ (72)
@@
$\widetilde{\bm{\Delta}}_{t}=\bm{\Delta}_{t}\mathds{1}(\mathcal{H}_{t}).$
@@
$\left\|\frac{1}{T}\sum_{i=1}^{T}\bm{Q}_{i}(\bm{A}_{i}-\bm{A})\widetilde{\bm{\Delta}}_{i-1}\right\|_{2}\leq 2\sqrt{2W_{\max}\log\frac{9}{\delta}},$
@@
$W_{\max}=\sup\frac{1}{T^{2}}\sum_{i=1}^{T}\|\bm{Q}_{i}(\bm{A}_{i}-\bm{A})\widetilde{\bm{\Delta}}_{i-1}\|_{2}^{2}$ $\lesssim\frac{1}{T^{2}}3\eta_{0}^{2}\left(\frac{2}{\beta}\right)^{\frac{1}{1-\alpha}}\cdot 4R^{2}\sum_{i=1}^{T}i^{-\alpha}$
@@
$\left\|\frac{1}{T}\sum_{i=1}^{T}\bm{Q}_{i}(\bm{A}_{i}-\bm{A})\widetilde{\bm{\Delta}}_{i-1}\right\|_{2}$ $\lesssim\eta_{0}\left(\frac{2}{\beta}\right)^{\frac{1}{2(1-\alpha)}}R\sqrt{\frac{1}{1-\alpha}\log\frac{9}{\delta}}$
@@
Combining (70), (71), (72), (73) and (B.2) with the triangle inequality and a union bound argument, we obtain with probability $1-\delta$ that
@@
$+2\sqrt{\frac{\mathsf{Tr}(\overline{\bm{\Lambda}}_{T})}{T}\log\frac{6d}{\delta}}+\frac{4\eta_{0}}{T}\left(\frac{2}{\beta}\right)^{\frac{1}{1-\alpha}}(2\|\bm{\theta}^{\star}\|_{2}+1)\log\frac{6d}{\delta}$

### v3

@@ A.1 Useful algebraic facts
$I_{2}$ $=\underset{I_{21}}{\underbrace{\frac{1}{T}\sum_{t=1}^{T}\bm{S}_{t}\bm{\Gamma}\bm{S}_{t}^{\top}}}+\underset{I_{22}}{\underbrace{\frac{1}{T}\sum_{t=1}^{T}\prod_{k=t}^{T}(\bm{I}-\eta_{k}\bm{A})\bm{\Lambda}^{\star}\prod_{k=t}^{T}(\bm{I}-\eta_{k}\bm{A}^{\top})}}$
($-\underset{I_{23}}{\underbrace{\left(\frac{1}{T}\sum_{t=1}^{T}\bm{S}_{t}\bm{\Gamma}\bm{A}^{-\top}\prod_{k=t}^{T}(\bm{I}-\eta_{k}\bm{A}^{\top})+\frac{1}{T}\sum_{t=1}^{T}\prod_{k=t}^{T}(\bm{I}-\eta_{k}\bm{A})\bm{A}^{-1}\bm{\Gamma}\bm{S}_{t}^{\top}\right)}}.$
Since $\|\bm{S}_{t}\|=O(T^{\alpha-1})$, it is easy to verify that $I_{21}=O(T^{2\alpha-2})\bm{\Gamma}$; meanwhile, theorem A.9 indicates that $I_{22}=T^{\alpha-1}\bm{X}(\bm{\Lambda}^{\star})+O(T^{2\alpha-2})\bm{\Lambda}^{\star}$. As for the term $I_{23}$, we observe that
$\left\|\sum_{t=1}^{T}\bm{S}_{t}\prod_{k=t}^{T}(\bm{I}-\eta_{k}\bm{A}^{\top})\right\|$ $\lesssim\widetilde{C}\sum_{t=1}^{T}t^{\alpha-1}\prod_{k=t}^{T}(1-\beta k^{-\alpha})$
$<\widetilde{C}T^{2\alpha-1}\sum_{t=1}^{T}t^{\alpha-1}\prod_{k=t+1}^{T}(1-\beta k^{-\alpha})$
$=\widetilde{C}T^{2\alpha-1}\cdot\frac{1}{\beta}\sum_{t=1}^{T}\left(\prod_{k=t+1}^{T}-\prod_{k=t}^{T}\right)(1-\beta k^{-\alpha})$
$<\widetilde{C}\frac{T^{2\alpha-1}}{\beta},$
where $\widetilde{C}$ is a problem-related quantity related to $\eta_{0},\alpha$ and $\gamma$. Consequently, it can be guaranteed that $I_{23}=O(T^{2\alpha-2})\bm{\Gamma}$. In combintation, we have shown that
$\overline{\bm{\Lambda}}_{T}-\bm{\Lambda}^{\star}$ $=I_{1}+I_{21}+I_{22}+I_{23}$
$=O(T^{-1})\bm{\Lambda}^{\star}+O(T^{2\alpha-2})\bm{\Gamma}+(T^{\alpha-1}\bm{X}(\bm{\Lambda}^{\star})+O(T^{2\alpha-2})\bm{\Lambda}^{\star})+O(T^{2\alpha-2})\bm{\Gamma}$
$=T^{\alpha-1}\bm{X}(\bm{\Lambda}^{\star})+O(T^{2\alpha-2})\bm{\Gamma}+O(T^{2\alpha-2})\bm{\Lambda}^{\star};$
the theorem follows naturally by combining this with the fact that $\bm{\Lambda}^{\star}=O(1)\bm{\Gamma}$, due to (31).∎
@@
With $\bm{A}$ defined as (4) and $\overline{\bm{\Lambda}}_{T}$ defined as in (49), when
@@ A.2.1 Berry-Esseen bounds
The following theorem, due to Shao and Zhang (2022), gives a Berry-Esseen bound for non-linear statistics which would be crucial for our analysis of the non-asymptotic distribution of the TD estimation error.
@@
(Theorem 2.1 of Shao and Zhang (2022)) Let $\bm{X}_{1},...,\bm{X}_{n}$ be independent random variables, and let $\bm{T}(\bm{X}_{1},...,\bm{X}_{n})$ be a $d$-dimensional statistic that can be represented as
@@ A.2.2 Properties of multivariate Gaussian distributions
The following theorem by Devroye et al. (2018) gives an upper bound for the total variation (TV) distance between two Gaussian random variables with same means and different covariance matrices.
Theorem A.14 (Devroye et al. (2018)).
Let $\bm{\Lambda}_{1},\bm{\Lambda}_{2}\in\mathbb{S}^{d\times d}$ be two positive definite matrices, and $\bm{\mu}$ be any vector in $\mathbb{R}^{d}$. Then the TV distance between Gaussian distributions $\mathcal{N}(\bm{\mu},\bm{\Lambda}_{1})$ and $\mathcal{N}(\bm{\mu},\bm{\Lambda}_{2})$ is bounded by
@@
In the case of $\bm{\Lambda}_{1}\succ\bm{\Lambda}_{2}$, the following proposition further specifies that the TV distance and the distance on convex sets are the same between $\mathcal{N}(\bm{0},\bm{\Lambda}_{1})$ and $\mathcal{N}(\bm{0},\bm{\Lambda}_{2})$.
@@
$\frac{p_{1}(\bm{x})}{p_{2}(\bm{x})}$ $=\frac{\frac{1}{\sqrt{\mathsf{det}(\bm{\Lambda_{1}})}}\exp(-\bm{x}^{\top}\bm{\Lambda_{1}}^{-1}\bm{x})}{\frac{1}{\sqrt{\mathsf{det}(\bm{\Lambda_{2}})}}\exp(-\bm{x}^{\top}\bm{\Lambda_{2}}^{-1}\bm{x})}$
@@
Combining these two inequalities completes the proof. ∎
@@ A.3 Concentration inequalities
The following Azuma’s inequality for matrix martingales is shown in Tropp (2011).
@@
Specifically, for vector-valued martingales, the following proposition is shown by Kallenberg and Sztencel (1991).
@@
The following Freedman’s inequality on Hilbert spaces was shown by Peng et al. (2024) in their Theorem 2.1.
Theorem A.19.
Let $\mathcal{X}$ be a Hilbert space, $\{\bm{X}_{i}\}_{i=1}^{n}$ be an $\mathcal{X}$-valued martingale difference sequence adapted to the filtration $\{\mathcal{F}_{i}\}_{i=1}^{n}$, $\bm{Y}_{i}:=\sum_{j=1}^{i}\bm{X}_{j}$ be the corresponding martingale, and $W_{i}=\sum_{j=1}^{i}\mathbb{E}\|\bm{X}_{j}\|^{2}\mid\mathscr{F}_{j-1}$. Suppose that $\max_{i\in[n]}\|\bm{X}_{i}\|\leq b$ for some constant $b>0$. Then for any $\varepsilon$ and $\sigma>0$, it can be guaranteed that
$\mathbb{P}\left(\exists k\in[n],\text{ s.t. }\|\bm{Y}_{k}\|\geq\varepsilon\text{ and }W_{k}\leq\sigma^{2}\right)\leq 2\exp\left(-\frac{\varepsilon^{2}/2}{\sigma^{2}+b\varepsilon/3}\right).$
@@ Appendix B Proof of main results
B.1 Proof of Theorem 3.1
In order to derive the high-probability convergence guarantee for $\overline{\bm{\Delta}}_{T}$, we firstly propose the following result on the time-uniform convergence of the original TD estimator $\bm{\Delta}_{t}$.
B.1.1 High-probability convergence guarantee for the original TD estimation error
@@
Consider the TD algorithm (7) with independent samples and decaying stepsizes $\eta_{t}=\eta_{0}t^{-\alpha}$, where $\alpha\in(\frac{1}{2},1)$. For any $\delta\in(0,1)$, there exists $\eta_{0}>0$, such that with probability at least $1-\delta$, the norm of $\bm{\Delta}_{t}$ is bounded by
@@
For simplicity, in this proof we denote
@@
For any given $\delta$, we define $t^{\star}$ as a function of $\delta$, specifically
@@
$\mathcal{H}_{i}=\{\max_{1\leq j\leq i}\|\bm{\Delta}_{j}\|\leq R\},\quad\text{and}\quad\tilde{\bm{\zeta}}_{i}=\bm{\zeta}_{i}\mathds{1}(\mathcal{H}_{i-1}).$
@@
$\|\tilde{\bm{\zeta}}_{i}\|_{2}\leq 4R+2\|\bm{\theta}^{\star}\|_{2}+1.$
@@
$\left\|\sum_{i=1}^{t}\bm{R}_{i}^{t}\tilde{\bm{\zeta}}_{i}\right\|_{2}^{2}$ $\leq\sum_{i=1}^{t}\|\bm{R}_{i}^{t}\|^{2}\|\tilde{\bm{\zeta}}_{i}\|_{2}^{2}$
@@
$\left\|\sum_{i=1}^{t}\bm{R}_{i}^{t}\tilde{\bm{\zeta}}_{i}\right\|_{2}$ $<2\eta_{0}(4R+2\|\bm{\theta}^{\star}\|_{2}+1)\sqrt{\frac{2}{2\alpha-1}\log\frac{9t^{\star}}{\delta}}$
@@
$=\left\|\prod_{k=1}^{t}(\bm{I}-\eta_{k}\bm{A})\bm{\Delta}_{0}\right\|_{2}+\left\|\sum_{i=1}^{t}\bm{R}_{i}^{t}\tilde{\bm{\zeta}}_{i}\right\|_{2}$
@@
$\left\|\sum_{i=1}^{t}\bm{R}_{i}^{t}\tilde{\bm{\zeta}}_{i}\right\|_{2}$ $\leq 2\eta_{0}(4R+2\|\bm{\theta}^{\star}\|_{2}+1)\sqrt{2W_{t}\log\frac{3}{\delta_{t}}}$
@@
$\left\|\sum_{i=1}^{t}\bm{R}_{i}^{t}\tilde{\bm{\zeta}}_{i}\right\|_{2}\leq 2\eta_{0}\left(\frac{2\alpha}{\beta}\right)^{\frac{\alpha}{2(1-\alpha)}}\sqrt{\frac{2}{2\alpha-1}\log\frac{9T}{\delta}}(4R+2\|\bm{\theta}^{\star}\|_{2}+1)t^{-\frac{\alpha}{2}}.$
@@
B.1.2 Completing the proof of Theorem 3.1
Similar to the proof of Theorem B.1, we firstly expand $\overline{\bm{\Delta}}_{T}$ into three terms: a non-random term capturing the error introduced by the intialization $\bm{\theta}_{0}$, a martingale, and a sum of $i.i.d.$ vectors. Specifically, (B.1.1) implies
@@
The term $I_{2}$ is a sum of $i.i.d.$ random vectors, and can hence be bounded by invoking the Freedman’s inequality on Hilbert spaces (Theorem A.19). Specifically, define
@@
$\left\|\frac{1}{T}\sum_{i=1}^{T}\bm{Q}_{i}(\bm{A}_{i}\bm{\theta}^{\star}-\bm{b}_{i})\right\|_{2}\leq 2\sqrt{W_{\max}\log\frac{2}{\delta}}+\frac{4}{3}B_{\max}\log\frac{2}{\delta}.$
@@
$\left\|\frac{1}{T}\sum_{i=1}^{T}\bm{Q}_{i}(\bm{A}_{i}\bm{\theta}^{\star}-\bm{b}_{i})\right\|_{2}\leq 2\sqrt{\frac{\mathsf{Tr}(\overline{\bm{\Lambda}}_{T})}{T}\log\frac{6}{\delta}}+\frac{4\eta_{0}}{T}\left(\frac{2}{\beta}\right)^{\frac{1}{1-\alpha}}(2\|\bm{\theta}^{\star}\|_{2}+1)\log\frac{6}{\delta}.$ (75)
@@
$\tilde{\bm{\Delta}}_{t}=\bm{\Delta}_{t}\mathds{1}(\mathcal{H}_{t}).$
@@
$\left\|\frac{1}{T}\sum_{i=1}^{T}\bm{Q}_{i}(\bm{A}_{i}-\bm{A})\tilde{\bm{\Delta}}_{i-1}\right\|_{2}\leq 2\sqrt{2W_{\max}\log\frac{9}{\delta}},$
@@
$W_{\max}=\sup\frac{1}{T^{2}}\sum_{i=1}^{T}\|\bm{Q}_{i}(\bm{A}_{i}-\bm{A})\tilde{\bm{\Delta}}_{i-1}\|_{2}^{2}$ $\lesssim\frac{1}{T^{2}}3\eta_{0}^{2}\left(\frac{2}{\beta}\right)^{\frac{1}{1-\alpha}}\cdot 4R^{2}\sum_{i=1}^{T}i^{-\alpha}$
@@
$\left\|\frac{1}{T}\sum_{i=1}^{T}\bm{Q}_{i}(\bm{A}_{i}-\bm{A})\tilde{\bm{\Delta}}_{i-1}\right\|_{2}$ $\lesssim\eta_{0}\left(\frac{2}{\beta}\right)^{\frac{1}{2(1-\alpha)}}R\sqrt{\frac{1}{1-\alpha}\log\frac{9}{\delta}}$
@@
Combining (73), (74), (75), (76) and (B.1.2) with the triangle inequality and a union bound argument, we obtain with probability $1-\delta$ that
@@
$+2\sqrt{\frac{\mathsf{Tr}(\overline{\bm{\Lambda}}_{T})}{T}\log\frac{6}{\delta}}+\frac{4\eta_{0}}{T}\left(\frac{2}{\beta}\right)^{\frac{1}{1-\alpha}}(2\|\bm{\theta}^{\star}\|_{2}+1)\log\frac{6}{\delta}$